Dolphins make sounds in air and water. What is the ratio of the wavelength of a sound in air to its wavelength in seawater if the speed of sound in water at that specific location is 1530 m/s and the air temperature is 25.00°C?
Answer #1

Speed of sound in air vs temperature
s = (331 + 0.606T) m/s

T is temperature in ºC

s= (331 + 0.606*25° C)
s = 346 m/s

air λ = s/f = 340/f
sea λ = s/f = 1530/f

ratio air/sea = 340/f / 1530/f = 340/1530 = 0.223

I hope it was helpful..